# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ C.2 Proof of Lemma 8
Figure 6: A schematic figure of the relations among Lemma 8, Lemma 32 and Lemma 33. The direction of the arrow indicates the convertibility with vanishing error. In Lemma 32, we find the growth rate of i.i.d. copies can be set to 1 by adjusting the scaling of the parameters $\bm{u}$ when analyzing the asymptotic convertibility. Lemma 33 shows that the asymptotic conversion between two pure-state models are possible if QGT is reduced during the conversion in the sense of matrix inequality. Combining these lemmas, Lemma 8 is proven.
@@
Let $\phi$ be an arbitrary pure state in a finite-dimensional Hilbert space. Fix a set of Hermitian operators $\bm{X}=\{X_{i}\}_{i=1}^{m}$. We define another set of Hermitian operator $\tilde{\bm{X}}=\{\tilde{X}_{i}\}_{i=1}^{m}$ by $\tilde{X}_{i}\coloneqq\Pi_{\phi}^{\perp}X_{i}\Pi_{\phi}+\Pi_{\phi}X_{i}\Pi_{\phi}^{\perp}$, where $\Pi_{\phi}\coloneqq\phi$ and $\Pi_{\phi}^{\perp}\coloneqq I-\phi$ are projectors onto a subspace spanned by $\phi$ and its orthogonal space, respectively. For $\epsilon\in(0,1/6)$, it holds
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\phi_{\bm{u}/\sqrt{N}}^{\otimes N},\tilde{\phi}_{\bm{u}/\sqrt{N}}^{\otimes N}\right)=0,$ (260)
where $\phi_{\bm{\theta}}\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}}\phi e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}}$ and $\tilde{\phi}_{\bm{\theta}}\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\tilde{\bm{X}}}\phi e^{-\mathrm{i}\bm{\theta}\cdot\tilde{\bm{X}}}$ with $\bm{\theta}\cdot\bm{X}=\sum_{i=1}^{m}\theta^{i}X_{i}$ and $\bm{\theta}\cdot\tilde{\bm{X}}=\sum_{i=1}^{m}\theta^{i}\tilde{X}_{i}$ for $\bm{\theta}\in\mathbb{R}^{m}$.
@@
Let us define $O\coloneqq\frac{1}{\sqrt{N}}\sum_{i=1}^{m}u^{i}X_{i}$ and $O^{\prime}\coloneqq\frac{1}{\sqrt{N}}\sum_{i=1}^{m}u^{i}\tilde{X}_{i}=\Pi_{\phi}^{\perp}O\Pi_{\phi}+\Pi_{\phi}O\Pi_{\phi}^{\perp}$ for $\bm{u}\in\mathbb{R}^{m}$ such that $\|\bm{u}\|<N^{\epsilon}$. Introducing a finite non-negative number $L$ by
$L\coloneqq\max\left\{\max_{\|\bm{\theta}\|=1}\|\bm{\theta}\cdot\bm{X}\|_{\infty},\max_{\|\bm{\theta}\|=1}\|\bm{\theta}\cdot\tilde{\bm{X}}\|_{\infty}\right\}$, (261)
we have $\|O\|_{\infty}\leq L\frac{\|\bm{u}\|}{\sqrt{N}}$ and $\|O^{\prime}\|_{\infty}\leq L\frac{\|\bm{u}\|}{\sqrt{N}}$. Applying Lemma 28 and Lemma 29 for $f(\theta)\coloneqq e^{-\mathrm{i}\theta O^{\prime}}e^{\mathrm{i}\theta O}\phi e^{-\mathrm{i}\theta O}e^{\mathrm{i}\theta O^{\prime}}$, we find
@@
$\leq\frac{2^{3}}{6}(\|O\|_{\infty}+\|O^{\prime}\|_{\infty})^{3}\leq\frac{2^{6}}{6}\left(L\frac{\|\bm{u}\|}{\sqrt{N}}\right)^{3}.$ (263)
@@
$\braket{\phi|O^{\prime n}O^{m}|\phi}=\begin{cases}1&\quad(n,m)=(0,0)\\ \braket{\phi|O|\phi}&\quad(n,m)=(0,1)\\ 0&\quad(n,m)=(1,0)\end{cases},$ (264)
@@
$\braket{\phi|O^{\prime n}O^{m}|\phi}=\begin{cases}\braket{\phi|O^{2}|\phi}&\quad(n,m)=(0,2)\\ \braket{\phi|O(I-\phi)O|\phi}&\quad(n,m)=(1,1),\,(2,0)\end{cases}$ (267)
@@
$\sum_{\begin{subarray}{c}n,m=0,1,2\\ n+m=2\end{subarray}}\frac{(-\mathrm{i})^{n}}{n!}\frac{\mathrm{i}^{m}}{m!}\braket{\phi|O^{\prime n}O^{m}|\phi}=-\frac{1}{2}\braket{\phi|O^{2}|\phi}+\frac{1}{2}\braket{\phi|O(I-\phi)O|\phi}=-\frac{1}{2}\braket{\phi|O|\phi}^{2}$ (268)
@@
$\braket{\phi|f^{\prime\prime}(0)|\phi}$ $=\sum_{\begin{subarray}{c}n,m,n^{\prime},m^{\prime}=0,1,2\\ n+m+n^{\prime}+m^{\prime}=2\end{subarray}}\frac{(-\mathrm{i})^{n}}{n!}\frac{\mathrm{i}^{m}}{m!}\frac{(-\mathrm{i})^{m^{\prime}}}{m^{\prime}!}\frac{\mathrm{i}^{n^{\prime}}}{n^{\prime}!}\braket{\phi|O^{\prime n}O^{m}|\phi}\braket{\phi|O^{m^{\prime}}O^{\prime n^{\prime}}|\phi}$ (269)
@@
Therefore, by using $\braket{\phi|f(1)|\phi}=\mathrm{Fid}\left(\phi_{\bm{u}/\sqrt{N}},\tilde{\phi}_{\bm{u}/\sqrt{N}}\right)$, Eq. (263) implies
$\left|\mathrm{Fid}\left(\phi_{\bm{u}/\sqrt{N}},\tilde{\phi}_{\bm{u}/\sqrt{N}}\right)-1\right|\leq\frac{2^{3}}{6}(\|O\|_{\infty}+\|O^{\prime}\|_{\infty})^{3}\leq\frac{2^{6}}{6}\left(L\frac{\|\bm{u}\|}{\sqrt{N}}\right)^{3}\leq\frac{2^{6}}{6}(LN^{-1/2+\epsilon})^{3}$ (271)
for any $\bm{u}$ such that $\|\bm{u}\|<N^{\epsilon}$.
Since $N^{-1/2+\epsilon}=o(1/N)$ if $\epsilon\in(0,1/6)$, we get
$\inf_{\|\bm{u}\|<N^{\epsilon}}\mathrm{Fid}\left(\phi_{\bm{u}/\sqrt{N}}^{\otimes N},\tilde{\phi}_{\bm{u}/\sqrt{N}}^{\otimes N}\right)=\inf_{\|\bm{u}\|<N^{\epsilon}}\left(\mathrm{Fid}\left(\phi_{\bm{u}/\sqrt{N}},\tilde{\phi}_{\bm{u}/\sqrt{N}}\right)\right)^{N}\geq\left(1+o\left(\frac{1}{N}\right)\right)^{N},$ (272)
@@
$\lim_{N\to\infty}\inf_{\|\bm{u}\|<N^{\epsilon}}\mathrm{Fid}\left(\phi_{\bm{u}/\sqrt{N}}^{\otimes N},\tilde{\phi}_{\bm{u}/\sqrt{N}}^{\otimes N}\right)$ $=\lim_{N\to\infty}\left(1+o\left(\frac{1}{N}\right)\right)^{N}=1.$ (273)
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\phi_{\bm{u}/\sqrt{N}}^{\otimes N},\tilde{\phi}_{\bm{u}/\sqrt{N}}^{\otimes N}\right)=0.$ (274)
Therefore, as long as the asymptotic properties of an i.i.d. copies of $\phi_{\bm{u}/\sqrt{N}}=e^{\mathrm{i}\bm{u}\cdot\bm{X}/\sqrt{N}}\phi e^{-\mathrm{i}\bm{u}\cdot\bm{X}/\sqrt{N}}$ are concerned, it suffice to analyze $\tilde{\phi}_{\bm{u}/\sqrt{N}}\coloneqq e^{\mathrm{i}\bm{u}\cdot\tilde{\bm{X}}/\sqrt{N}}\phi e^{\mathrm{i}\bm{u}\cdot\tilde{\bm{X}}/\sqrt{N}}$.
The asymptotic behavior of $\tilde{\phi}_{\bm{u}/\sqrt{N}}^{\otimes N}$ can be understood by studies on QLAN. The result in [87] is complicated since it aims to prove a general statement applicable to a wide class of statistical models. In this paper, we only focus on pure-state models. For pure-state unitary models, the result in [87] can be stated as follows:
@@
[…]
We introduce a $(d-1)\times m$ matrix defined by
$\mathrm{i}C_{ki}\coloneqq\braket{k|\tilde{X}_{i}|0},$ (276)
[…]
Let $\mathcal{H}_{d-1}$ denote the (d1)(d-1)-mode Fock space. For $\bm{z}=(z_{1},\cdots,z_{d-1})^{\top}\in\mathbb{C}^{d-1}$, we denote the coherent state by
@@
[…]
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{E}_{N}\left(\phi_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N}\right),\phi_{\bm{u}/\sqrt{N}}^{\otimes\lfloor rN\rfloor}\right)$ $=0,$ (280)
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\phi_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N},\mathcal{E}_{N}^{\prime}\left(\phi_{\bm{u}/\sqrt{N}}^{\otimes\lfloor rN\rfloor}\right)\right)$ $=0$ (281)
@@
For $\tilde{X}_{i}\coloneqq(I-\phi)X_{i}\phi+\phi X_{i}(I-\phi)$, consider a sequence of statistical models $\{(\tilde{\phi}_{\bm{\theta}})^{\otimes\lfloor rN\rfloor}\}_{N}$, where $r>0$ and $\tilde{\phi}_{\bm{\theta}}\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\tilde{\bm{X}}}\phi e^{-\mathrm{i}\bm{\theta}\cdot\tilde{\bm{X}}}$. From Lemma 31, there exist sequences of quantum channels $\{\mathcal{T}_{N}\}_{N}$ and $\{\mathcal{S}_{N}\}_{N}$ such that
@@
for $\epsilon\in(0,1/9)$. Introducing a new parameter $\bm{u}^{\prime}/\sqrt{N}\coloneqq\bm{u}/\sqrt{\lfloor rN\rfloor}$, we get
@@
$\lim_{N\to\infty}\left|\Braket{\sqrt{\frac{\lfloor rN\rfloor}{N}}C\bm{u}^{\prime}|\sqrt{r}C\bm{u}^{\prime}}\right|=\lim_{N\to\infty}e^{-\|C\bm{u}^{\prime}\|^{2}\left(\sqrt{r}-\sqrt{\frac{\lfloor rN\rfloor}{N}}\right)^{2}}=1$ (287)
for any $\bm{u}^{\prime}\in\mathbb{R}^{m}$. Therefore, we get
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{S}_{\lfloor rN\rfloor}\circ\mathcal{T}_{N}\left(\tilde{\phi}_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N}\right),\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}\right)$ $=0,$ (292)
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\tilde{\phi}_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N},\mathcal{S}_{N}\circ\mathcal{T}_{\lfloor rN\rfloor}\left(\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}\right)\right)$ $=0$ (293)
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{S}_{\lfloor rN\rfloor}\circ\mathcal{T}_{N}\left(\phi_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N}\right),\phi^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}\right)$ $=0,$ (294)
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\phi_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N},\mathcal{S}_{N}\circ\mathcal{T}_{\lfloor rN\rfloor}\left(\phi^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}\right)\right)$ $=0$ (295)
@@
Importantly, for both $\phi_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N}$ and $\phi^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}$, the QGT with respect to the parameters $\bm{u}$ is equal to $r\,\mathrm{Cov}(\phi,\bm{X})$ in the limit of $N\to\infty$, where the elements of covariance matrix are defined by $\left(\mathrm{Cov}(\phi,\bm{X})\right)_{ij}\coloneqq\braket{\phi|X_{i}(I-\phi)X_{j}|\phi}$. In other words, the QGTs are preserved under reversible asymptotic conversion between $\{\phi_{\sqrt{r}\bm{u}/\sqrt{N}}^{\otimes N}\}_{N}$ and $\{\phi^{\otimes\lfloor rN\rfloor}_{\bm{u}/\sqrt{N}}\}_{N}$.

### v4

@@ D.2.2 Proof of Lemma 25
Figure 4: A schematic figure of the relations among Lemma 25, Lemma 32 and Lemma 33. The direction of the arrow indicates the convertibility with vanishing error. In Lemma 32, we find the growth rate of i.i.d. copies can be set to 1 by adjusting the scaling of the parameters $u$ when analyzing the asymptotic convertibility. Lemma 33 shows that the asymptotic conversion between two pure-state models are possible if QGT is reduced during the conversion in the sense of matrix inequality. Combining these lemmas, Lemma 25 is proven.
@@
Let $\phi$ be an arbitrary pure state in a finite-dimensional Hilbert space. Fix a set of Hermitian operators $\{X_{\mu}^{\prime}\}_{\mu=1}^{m}$. We define another set of Hermitian operator $\{\tilde{X}_{\mu}^{\prime}\}_{\mu=1}^{m}$ by $\tilde{X}_{\mu}^{\prime}\coloneqq(I-\phi)X_{\mu}^{\prime}\phi+\phi X_{\mu}^{\prime}(I-\phi)$. For $\epsilon\in(0,1/6)$, it holds
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\phi_{u/\sqrt{N}}^{\otimes N},\tilde{\phi}_{u/\sqrt{N}}^{\otimes N}\right)=0,$ (263)
where $\phi_{\theta}\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}^{\prime}}\phi e^{-\mathrm{i}\theta^{\mu}X_{\mu}^{\prime}}$ and $\tilde{\phi}_{\theta}\coloneqq e^{\mathrm{i}\theta^{\mu}\tilde{X}_{\mu}^{\prime}}\phi e^{-\mathrm{i}\theta^{\mu}\tilde{X}_{\mu}^{\prime}}$ for $\theta\in\mathbb{R}^{m}$.
@@
Let us define $O\coloneqq\frac{1}{\sqrt{N}}u^{\mu}X_{\mu}^{\prime}$ and $\tilde{O}\coloneqq\frac{1}{\sqrt{N}}u^{\mu}\tilde{X}_{\mu}^{\prime}=(I-\phi)O\phi+\phi O(I-\phi)$ for $u\in\mathbb{R}^{m}$ such that $\|u\|<N^{\epsilon}$. Introducing a finite non-negative number $L$ by
$L\coloneqq\max\left\{\max_{\|\theta\|=1}\|\theta^{\mu}X_{\mu}^{\prime}\|_{\infty},\max_{\|\theta\|=1}\|\theta^{\mu}\tilde{X}_{\mu}^{\prime}\|_{\infty}\right\},$ (264)
we have $\|O\|_{\infty}\leq L\frac{\|u\|}{\sqrt{N}}$ and $\|\tilde{O}\|_{\infty}\leq L\frac{\|u\|}{\sqrt{N}}$. Applying Lemma 28 and Lemma 29 for $f(\theta)\coloneqq e^{-\mathrm{i}\theta\tilde{O}}e^{\mathrm{i}\theta O}\phi e^{-\mathrm{i}\theta O}e^{\mathrm{i}\theta\tilde{O}}$, we find
@@
$\leq\frac{2^{3}}{6}(\|O\|_{\infty}+\|\tilde{O}\|_{\infty})^{3}\leq\frac{2^{6}}{6}\left(L\frac{\|u\|}{\sqrt{N}}\right)^{3}.$ (266)
@@
$\braket{\phi|\tilde{O}^{n}O^{m}|\phi}=\begin{cases}1&\quad(n,m)=(0,0)\\ \braket{\phi|O|\phi}&\quad(n,m)=(0,1)\\ 0&\quad(n,m)=(1,0)\end{cases},$ (267)
@@
$\braket{\phi|\tilde{O}^{n}O^{m}|\phi}=\begin{cases}\braket{\phi|O^{2}|\phi}&\quad(n,m)=(0,2)\\ \braket{\phi|O(I-\phi)O|\phi}&\quad(n,m)=(1,1),\,(2,0)\end{cases}$ (270)
@@
$\sum_{\begin{subarray}{c}n,m=0,1,2\\ n+m=2\end{subarray}}\frac{(-\mathrm{i})^{n}}{n!}\frac{\mathrm{i}^{m}}{m!}\braket{\phi|\tilde{O}^{n}O^{m}|\phi}=-\frac{1}{2}\braket{\phi|O^{2}|\phi}+\frac{1}{2}\braket{\phi|O(I-\phi)O|\phi}=-\frac{1}{2}\braket{\phi|O|\phi}^{2}$ (271)
@@
$\braket{\phi|f^{\prime\prime}(0)|\phi}$ $=\sum_{\begin{subarray}{c}n,m,n^{\prime},m^{\prime}=0,1,2\\ n+m+n^{\prime}+m^{\prime}=2\end{subarray}}\frac{(-\mathrm{i})^{n}}{n!}\frac{\mathrm{i}^{m}}{m!}\frac{(-\mathrm{i})^{m^{\prime}}}{m^{\prime}!}\frac{\mathrm{i}^{n^{\prime}}}{n^{\prime}!}\braket{\phi|\tilde{O}^{n}O^{m}|\phi}\braket{\phi|O^{m^{\prime}}\tilde{O}^{n^{\prime}}|\phi}$ (272)
@@
Therefore, by using $\braket{\phi|f(1)|\phi}=\mathrm{Fid}\left(\phi_{u/\sqrt{N}},\tilde{\phi}_{u/\sqrt{N}}\right)$, Eq. (266) implies
$\left|\mathrm{Fid}\left(\phi_{u/\sqrt{N}},\tilde{\phi}_{u/\sqrt{N}}\right)-1\right|\leq\frac{2^{3}}{6}(\|O\|_{\infty}+\|\tilde{O}\|_{\infty})^{3}\leq\frac{2^{6}}{6}\left(L\frac{\|u\|}{\sqrt{N}}\right)^{3}\leq\frac{2^{6}}{6}(LN^{-1/2+\epsilon})^{3}$ (274)
for any $u$ such that $\|u\|<N^{\epsilon}$.
Since $(N^{-1/2+\epsilon})^{3}=o(1/N)$ if $\epsilon\in(0,1/6)$, we get
$\inf_{\|u\|<N^{\epsilon}}\mathrm{Fid}\left(\phi_{u/\sqrt{N}}^{\otimes N},\tilde{\phi}_{u/\sqrt{N}}^{\otimes N}\right)=\inf_{\|u\|<N^{\epsilon}}\left(\mathrm{Fid}\left(\phi_{u/\sqrt{N}},\tilde{\phi}_{u/\sqrt{N}}\right)\right)^{N}\geq\left(1+o\left(\frac{1}{N}\right)\right)^{N},$ (275)
@@
$\lim_{N\to\infty}\inf_{\|u\|<N^{\epsilon}}\mathrm{Fid}\left(\phi_{u/\sqrt{N}}^{\otimes N},\tilde{\phi}_{u/\sqrt{N}}^{\otimes N}\right)$ $=\lim_{N\to\infty}\left(1+o\left(\frac{1}{N}\right)\right)^{N}=1.$ (276)
@@
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\phi_{u/\sqrt{N}}^{\otimes N},\tilde{\phi}_{u/\sqrt{N}}^{\otimes N}\right)=0.$ (277)
Therefore, as long as the asymptotic properties of an i.i.d. copies of $\phi_{u/\sqrt{N}}=e^{\mathrm{i}u^{\mu}X_{\mu}^{\prime}/\sqrt{N}}\phi e^{-\mathrm{i}u^{\nu}X_{\nu}^{\prime}/\sqrt{N}}$ are concerned, it suffice to analyze $\tilde{\phi}_{u/\sqrt{N}}=e^{\mathrm{i}u^{\mu}\tilde{X}_{\mu}^{\prime}/\sqrt{N}}\phi e^{-\mathrm{i}u^{\nu}\tilde{X}_{\nu}^{\prime}/\sqrt{N}}$.
The asymptotic behavior of $\tilde{\phi}_{u/\sqrt{N}}^{\otimes N}$ can be understood by studies on QLAN. The result in [73] is complicated since it aims to prove a general statement applicable to a wide class of statistical models. In this paper, we only focus on pure-state models. For pure-state unitary models, the result in [73] can be stated as follows:
@@
Let $\phi$ be an arbitrary state in a $d$-dimensional Hilbert space. For any fixed set of Hermitian operators $\{X_{\mu}^{\prime}\}_{\mu=1}^{m}$, we define a pure-state qudit model $\tilde{\phi}_{\theta}\coloneqq e^{\mathrm{i}\theta^{\mu}\tilde{X}_{\mu}^{\prime}}\phi e^{-\mathrm{i}\theta^{\nu}\tilde{X}_{\nu}^{\prime}}$ for $\theta\in\mathbb{R}^{m}$ with $\tilde{X}_{\mu}^{\prime}\coloneqq(I-\phi)X_{\mu}^{\prime}\phi+\phi X_{\mu}^{\prime}(I-\phi)$.
@@
$\mathrm{i}C_{k\mu}^{\prime}\coloneqq\braket{k|\tilde{X}_{\mu}^{\prime}|\phi}$ (278)
for $k=1,\ldots,d-1$ and $\mu=1,\ldots,m$.
Let $\mathcal{H}_{d-1}$ denote the (d1)(d-1)-mode Fock space. For $z=(z_{1},\cdots,z_{d-1})^{\top}\in\mathbb{C}^{d-1}$, we denote the coherent state by
@@
Let $\phi$ be an arbitrary pure state in a finite-dimensional Hilbert space. Fix any sets of Hermitian operators $\{X_{\mu}^{\prime}\}_{\mu=1}^{m}$ and introduce a pure-state statistical model $\phi_{\theta}\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}^{\prime}}\phi e^{-\mathrm{i}\theta^{\nu}X_{\nu}^{\prime}}$ for $\theta\in\mathbb{R}^{m}$. For any $r>0$, there exist sequences of quantum channels $\{\mathcal{E}_{N}\}_{N}$ and $\{\mathcal{E}^{\prime}_{N}\}_{N}$ such that
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}_{N}\left(\phi_{\sqrt{r}u/\sqrt{N}}^{\otimes N}\right),\phi_{u/\sqrt{N}}^{\otimes\lfloor rN\rfloor}\right)$ $=0,$ (282)
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\phi_{\sqrt{r}u/\sqrt{N}}^{\otimes N},\mathcal{E}_{N}^{\prime}\left(\phi_{u/\sqrt{N}}^{\otimes\lfloor rN\rfloor}\right)\right)$ $=0$ (283)
@@
For $\tilde{X}_{\mu}^{\prime}\coloneqq(I-\phi)X_{\mu}^{\prime}\phi+\phi X_{\mu}^{\prime}(I-\phi)$, consider a sequence of statistical models $\{(\tilde{\phi}_{\theta})^{\otimes\lfloor rN\rfloor}\}_{N}$, where $r>0$ and $\tilde{\phi}_{\theta}\coloneqq e^{\mathrm{i}\theta^{\mu}\tilde{X}_{\mu}^{\prime}}\phi e^{-\mathrm{i}\theta^{\nu}\tilde{X}_{\nu}^{\prime}}$. From Lemma 31, there exist sequences of quantum channels $\{\mathcal{T}_{N}\}_{N}$ and $\{\mathcal{S}_{N}\}_{N}$ such that
@@
for $\epsilon\in(0,1/9)$. Introducing a new parameter $u^{\prime}/\sqrt{N}\coloneqq u/\sqrt{\lfloor rN\rfloor}$, we get
@@
$\lim_{N\to\infty}\left|\Braket{\sqrt{\frac{\lfloor rN\rfloor}{N}}C^{\prime}u^{\prime}|\sqrt{r}C^{\prime}u^{\prime}}\right|=\lim_{N\to\infty}e^{-\|C^{\prime}u^{\prime}\|^{2}\left(\sqrt{r}-\sqrt{\frac{\lfloor rN\rfloor}{N}}\right)^{2}}=1$ (289)
for any $u^{\prime}\in\mathbb{R}^{m}$. Therefore, we get
@@
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{S}_{\lfloor rN\rfloor}\circ\mathcal{T}_{N}\left(\tilde{\phi}_{\sqrt{r}u/\sqrt{N}}^{\otimes N}\right),\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{u/\sqrt{N}}\right)$ $=0,$ (294)
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\tilde{\phi}_{\sqrt{r}u/\sqrt{N}}^{\otimes N},\mathcal{S}_{N}\circ\mathcal{T}_{\lfloor rN\rfloor}\left(\tilde{\phi}^{\otimes\lfloor rN\rfloor}_{u/\sqrt{N}}\right)\right)$ $=0$ (295)
@@
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{S}_{\lfloor rN\rfloor}\circ\mathcal{T}_{N}\left(\phi_{\sqrt{r}u/\sqrt{N}}^{\otimes N}\right),\phi^{\otimes\lfloor rN\rfloor}_{u/\sqrt{N}}\right)$ $=0,$ (296)
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\phi_{\sqrt{r}u/\sqrt{N}}^{\otimes N},\mathcal{S}_{N}\circ\mathcal{T}_{\lfloor rN\rfloor}\left(\phi^{\otimes\lfloor rN\rfloor}_{u/\sqrt{N}}\right)\right)$ $=0$ (297)
@@
Importantly, for both $\phi_{\sqrt{r}u/\sqrt{N}}^{\otimes N}$ and $\phi^{\otimes\lfloor rN\rfloor}_{u/\sqrt{N}}$, the QGT with respect to the parameters $u$ is equal to $r\,\mathrm{Cov}(\phi,\{X_{\mu}^{\prime}\})$ in the limit of $N\to\infty$, where the elements of covariance matrix are defined by $\left(\mathrm{Cov}(\phi,\{X_{\mu}^{\prime}\})\right)_{\mu\nu}\coloneqq\mathrm{Tr}(\phi X_{\mu}^{\prime}(I-\phi)X_{\nu}^{\prime})$. In other words, the QGTs are preserved under reversible asymptotic conversion between $\{\phi_{\sqrt{r}u/\sqrt{N}}^{\otimes N}\}_{N}$ and $\{\phi^{\otimes\lfloor rN\rfloor}_{u/\sqrt{N}}\}_{N}$.
@@
Let $\psi\in\mathcal{P}(\mathcal{H})$ and $\phi\in\mathcal{P}(\mathcal{H}^{\prime})$ be arbitrary pure states in finite-dimensional Hilbert spaces $\mathcal{H},\mathcal{H}^{\prime}$. Fix any sets of Hermitian operators $\{X_{\mu}\}_{\mu=1}^{m}$ and $\{X_{\mu}^{\prime}\}_{\mu=1}^{m}$. If $\mathrm{Cov}(\psi,\{X_{\mu}\}_{\mu=1}^{m})\geq\mathrm{Cov}(\phi,\{X_{\mu}^{\prime}\}_{\mu=1}^{m})$ holds, then there exists a quantum channel $\mathcal{E}$ such that
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}^{\otimes N}\left(\psi_{u/\sqrt{N}}^{\otimes N}\right),\phi_{u/\sqrt{N}}^{\otimes N}\right)=0$ (298)
for $\epsilon\in(0,1/6)$, where $\psi_{\theta}\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}}\psi e^{-\mathrm{i}\theta^{\nu}X_{\nu}}$ and $\phi_{\theta}\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}^{\prime}}\phi e^{-\mathrm{i}\theta^{\nu}X_{\nu}^{\prime}}$ for $\theta\in\mathbb{R}^{m}$.
Proof.
Following Lemma 30, we introduce $\tilde{X}_{\mu}\coloneqq(I-\psi)X_{\mu}\psi+\psi X_{\mu}(I-\psi)$ and $\tilde{X}_{\mu}^{\prime}\coloneqq(I-\phi)X_{\mu}^{\prime}\phi+\phi X_{\mu}^{\prime}(I-\phi)$ and denote
$\tilde{\psi}_{u/\sqrt{N}}\coloneqq e^{\mathrm{i}\frac{1}{\sqrt{N}}u^{\mu}\tilde{X}_{\mu}}\psi e^{-\mathrm{i}\frac{1}{\sqrt{N}}u^{\nu}\tilde{X}_{\nu}},\quad\tilde{\phi}_{u/\sqrt{N}}\coloneqq e^{\mathrm{i}\frac{1}{\sqrt{N}}u^{\mu}\tilde{X}_{\mu}^{\prime}}\phi e^{-\mathrm{i}\frac{1}{\sqrt{N}}u^{\nu}\tilde{X}_{\nu}^{\prime}}.$ (299)